# Rotation Control Unlearning: Quantifying and Controlling Continuous Unlearning for LLM with The Cognitive Rotation Space

### v1

@@ 1 Introduction
However, unlearning methods in LLM are often not a one-time operation but a continuous process in real world. Most of them exist the cumulative catastrophic utility loss (Gao et al. (2025)) when dealing with continuous unlearning. Furthermore, they still require a retained dataset to maintain the model’s utility. This retained dataset consists of a part of the original training dataset (Bourtoule et al. (2021)). Since LLM require a large amount of data for training (Wang et al. (2024)), using the retained dataset in continuous unlearning is not feasible (Liu et al. (2025)). The work of $o^{3}$ (Gao et al. (2025)) proposed an unlearning paradigm for allocating weights to LoRA (Hu et al. (2022)) based on the Out-Of-Distribution (OOD) detector score to address the aforementioned issues. However, its main limitation lies in the lack of in-depth investigation into the continuous unlearning process, which prevents $o^{3}$ from effectively quantifying the unlearning process. Moreover, the mapping relationship between Out-Of-Distribution (OOD) scores and LoRA weights still requires extensive experimentation to determine, resulting in a lack of precise control over the unlearning process.
In this work, we propose Rotation Control Unlearning (RCU), a novel unlearning method that addresses the above challenges by conceptualizing LoRA update as rotations within a cognitive rotation space. Our approach introduces a skew symmetric loss in LoRA update paradigm to formulate the unlearning process as rotation operations, with the rotational angle serving as a precise quantification metric. We introduce an orthogonal rotation axes loss to enforce perpendicular rotation directions for consecutive unlearning requests, effectively mitigating cumulative catastrophic utility loss by minimizing inter unlearning request interference. Furthermore, to enhance compatibility, we design an unlearning alignment loss that guides the OOD detector to produce representations aligned with our LoRA update paradigm. These representations then collaborate with the distributional shift compensator to generate rotational salience weights for auxiliary quantification.
@@ 2 Preliminary
Figure 1: The overall architecture of our method is shown in the figure. In the training pipeline, the orthogonal rotation axes loss $\mathcal{L}_{o}$ is applied to the attention layers of the LLMs for training; simultaneously, the unlearning alignment loss $\mathcal{L}_{Ua}$ is used to train an OOD detector, whose output is fed into the distributional shift compensator to generate the rotational salience weight $\beta$. In the inference pipeline, given that the LoRA parameters BABA are proportional to the rotation angle $\theta_{R_{BA}}$ in the Cognitive Rotation Space $R_{BA}$. We control the rotation angle $\theta_{R_{BA}}$ amplitude by adjusting the scale of LoRA BABA, and use the weight $\beta$ to dynamically load the parameters that match the required unlearning degree.
@@ 3.1 Continuous Unlearning for LLM with LoRA
The continuous unlearning process of LLM inevitably leads to cumulative catastrophic utility loss (Gao et al. (2025)). This requires the method to simultaneously achieve the continuous preservation of its historical unlearning knowledge and the original utility when handling the current unlearning request. In addition, the unlearning methods often require a lot of computational resources and cost. We proposed the RCU to address these challenges. Within this framework, we constructed the cognitive rotation space through a new LoRA update paradigm and regarded the parameter updates of LoRA as rotations within this space. By monitoring and manipulating the changes in this rotation angle $\theta$, we were able to achieve controllable adjustment of the degree of unlearning LoRA loading.
@@
$W\leftarrow I+BA.$ (3)
In order to better control the unlearning process of LoRA. Our method constructs a new LoRA update paradigm from the perspective of rotation angle.
@@
However, every element $C$ in the Lie algebra $\mathfrak{so}\left(n\right)$ can be mapped to the $R$ in the Lie group $\mathit{SO}\left(n\right)$ by the exponential map $exp\left(C\right)$. From this we obtain:
@@
In addition, due to the influence of factors such as the learning rate, $BA<<I$ (the parameter size information of $BA$ is provided in the appendix.). Thereby, there exists a cognitive rotation space $R_{BA}\approx I+BA$.
@@
The orthogonal rotation axes loss are as follows:
$\mathcal{L}_{o}=\left\|\left(I+W_{t}-W_{t-1}\right)\cdot W_{t-1}\right\|_{F}^{2}=\left\|\left(I+B_{t}A_{t}-B_{t-1}A_{t-1}\right)\cdot\left(B_{t-1}A_{t-1}\right)\right\|_{F}^{2},$ (11)
@@
here,we set $\lambda_{1}=0.1$$\lambda_{2}=0.1$ and $\lambda_{3}=1$ in ScienceQA dataset. We set $\lambda_{1}=0.01$$\lambda_{2}=0.5$ and $\lambda_{3}=1$ in TOFU dataset.
@@ 4.2 Experimental Setup
Table 1: Performance Comparison between our method and other baselines when continually unlearning TOFU-forget01, -forget05, and -forget10 in Fictitious Knowledge Generation.
@@ 4.3 Experimental Results
Question Answering. The results of our method are shown in Figure 2 (c)(d) and Figure 3. We compared base, PO, SOPO, $o^{3}$ and our results. Some detailed results can be found in the appendix. Our method achieved the same results as the base method (C.QA. and O.QA.). The results of the R.D. were also very similar to those of the base method. However, our method performed much better than the others in the S.U. and D.U. indicators. Compared with the most advanced method $o^{3}$, the S.U. decreased by 9.68%, 4.02%, 4.05%, 7.58% and 10.82% respectively on five requests. The D.U. compared to the most advanced method $o^{3}$, on average, decreased by 17.67% across five requests. In addition, the number of training parameters of our method is much smaller than that of $o^{3}$.
Fictitious Knowledge Generation. Table 1 presents the experimental results of our method on the TOFU dataset. Based on the results, we found that our method can effectively enhance the LLMs’ ability to unlearning (D.U. and S.U.). Compared with the currently best continuous unlearning method $o^{3}$, the result on the S.U. decreased by 6.23%, 20.6% and 10.3%. the result on the D.U. decreased by 12.5%, 17.5% and 7.5%. Furthermore, in terms of the stability of the LLMs’ utility, the performance level of our method is comparable to the current best results. We observed that as the number of unlearning requests increased, the unlearning accuracy on the test set (D.U.) of most methods showed significant fluctuations or increasing. This phenomenon indicates that catastrophic utility loss continues to accumulate, and the LLMs’ stability is affected. In contrast, our method not only effectively mitigates such utility losses but also significantly improves the performance stability of LLMs during the continuous forgetting process.
@@
Figure 2: Experimental results on the ScienceQA dataset. (a)The relationship between the $\beta$ and unlearning processes (S.U.). (b) The relationship between the $\beta$ and unlearning processes (D.U.). (c) The results of C.QA.. (d)The results of O.QA..
4.4 Hyperparameter Analysis
In Figure 2 (a) and Figure 2 (b), we have presented the changes about $\beta$, along with the experimental results for each unlearning request. Based on the results in ScienceQA dataset, we can observe that when $\beta$ is within the range of 0.15 to 0.45, the model rapidly undergoes the process of unlearning. However, when $\beta$ is less than 0.15, the model does not forget the knowledge. And when $\beta$ is greater than 0.45, the model has achieved complete forgetting and no longer continues to unlearning. The above experimental results indicate that the unlearning process of RCU is achieved within a very small rotation angle change range in the cognitive rotation space. Therefore, when designing the OOD detector and the distributional shift compensator, we must also ensure that the rotational salience weights they output have a relatively concentrated distribution to match the above characteristics. The $\beta$ changes regarding TOFU dataset are presented in the appendix.
@@ 5 Conclusion
Existing LLM machine unlearning methods rely on a retained dataset to maintain model utility, which makes them susceptible to cumulative catastrophic utility loss when handling continuous unlearning requests. To overcome this limitation, we introduce RCU, a retain-free approach that formulates LoRA updates as rotations within a specially constructed cognitive rotation space. This formulation allows the degree of unlearning to be directly quantified through corresponding changes in rotational angles. Furthermore, by analyzing the update dynamics during unlearning, we propose a rotational salience weight to achieve precise and continuous control over the unlearning process. Our method is supported by theoretical guarantees, and we rigorously establish its efficacy through a mathematical analysis. Extensive experiments demonstrate that RCU achieves superior unlearning effectiveness while maintaining model utility on multiple datasets.
@@ Appendix A Appendix
• A.4: An in-depth study on why our method is effective.
• A.5: The experimental results on the TOFU dataset showing the relationship between $\beta$ and the unlearning process.
• A.6: The detailed calculation steps of the OOD Score.
@@
$exp\left(kC\right)$ $=exp\left(C^{\prime}\right)=Q\cdot exp\left(diag\left(0,...,0,B\left(k\theta_{1}\right),...,B\left(k\theta_{m}\right)\right)\right)\cdot Q^{T}$ (22)
$=Q\cdot diag\left(1,...,1.B\left(k\theta_{1}\right),...,B\left(k\theta_{m}\right)\right)\cdot Q^{T}.$ (23)
@@ A.2 More Experimental Results
The detailed results of our work on the ScienceQA dataset are shown in Table 3. Based on the results, we can observe that compared to $o^{3}$, our method is more stable and has a better ability to resist cumulative catastrophic utility loss.
Table 3: Performance Comparison between our method and other baselines when continually unlearning biology, physics, chemistry, economics and earth-science in Fictitious Knowledge Generation. The unlearning effectiveness is measured by the generation accuracy of the unlearning train data and unlearning test data denoted as S.U. and D.U., CommonsenseQA (C.QA.), OpenbookQA (O.QA.) respectively. Utility preservation is evaluated by the generation accuracy of Retained Distribution (R.D.).
@@ A.3 Large Language Model Usage Declaration
A.4 Research on Our Method
Unlearning Request Param $B$ Param AA
biology $10^{-6}$ $10^{-5}$
physice $10^{-5}$ $10^{-5}$
chemistry $10^{-5}$ $10^{-5}$
economics $10^{-6}$ $10^{-5}$
earth-science $10^{-6}$ $10^{-5}$
A.5 Hyperparameter analysis on the TOFU dataset

### v3

@@ 1 Introduction
However, unlearning methods in LLM are often not a one-time operation but a continuous process in real world. Most of them exist the cumulative catastrophic utility loss (Gao et al. (2025)) when dealing with continuous unlearning. The cumulative catastrophic utility loss causes a significant decline in both the LLM’s unlearning capability and utility retention capacity during the continuous unlearning process as the number of requests increases. At the parameter level, this manifests as new unlearning requests inducing parameter shift in previously learned ones. Furthermore, they still require a retained dataset to maintain the model’s utility. This retained dataset consists of a part of the original training dataset (Bourtoule et al. (2021)). Since LLM require a large amount of data for training (Wang et al. (2024)), using the retained dataset in continuous unlearning is not feasible (Liu et al. (2025)).
The work of $o^{3}$ (Gao et al. (2025)) proposes to mitigate cumulative catastrophic utility loss by imposing orthogonal constraints on LoRA parameters and introduces weights for the LoRA (Hu et al. (2022)) modules to represent the degree of unlearning. However, this approach suffers from several significant limitations. Firstly, the effectiveness of its simple orthogonal constraints on LoRA parameters diminishes as the number of unlearning requests increases, making it difficult to sustainably alleviate cumulative catastrophic utility loss. Secondly, using LoRA weights to quantify the degree of unlearning lacks interpretable justification. Finally, the mapping from the Out-Of-Distribution (OOD) detector outputs to the corresponding weights heavily relies on empirical design, which substantially increases the complexity and cost of application.
In this work, we propose Rotation Control Unlearning (RCU), a novel unlearning method that addresses the above challenges. This method is inspired by the theory of Lie group (Gallier (2001)) and re-constructs the LoRA update paradigm through mathematical derivation. It re-expresses the unlearning update of LLM as rotational operations within a cognitive rotation space. The cognitive rotation space is defined as a high-dimensional rotation space, which is used to depict the rotational transformations experienced by the original parameters of LLM during continuous unlearning. This enables the transformation of the uncontrollable parameter shift into controllable rotational angle changes, thereby effectively alleviating the cumulative catastrophic utility loss. The specific mathematical formulation is elaborated in the methodology. Our approach introduces a skew symmetric loss in LoRA update paradigm to formulate the unlearning process as rotation operations, with the rotational angle serving as a precise quantification metric. We introduce an orthogonal rotation axes loss to enforce perpendicular rotation directions for consecutive unlearning requests, effectively mitigating cumulative catastrophic utility loss by minimizing inter unlearning request interference. Furthermore, to enhance compatibility, we design an unlearning alignment loss that guides the OOD detector to produce representations aligned with our LoRA update paradigm. These representations then collaborate with the distributional shift compensator to generate rotational salience weights for auxiliary quantification. Finally, our method is supported by straightforward experimental interpretability and requires significantly fewer trainable parameters than $o^{3}$.
@@
  • We demonstrate the connection between rotation and unlearning through mathematical proof and experimental validation.
@@ 2 Preliminary
Figure 1: The overall architecture of our method is shown in the figure. In the training pipeline, the orthogonal rotation axes loss $\mathcal{L}_{o}$ is applied to the attention layers of the LLMs for training; simultaneously, the unlearning alignment loss $\mathcal{L}_{Ua}$ is used to train an OOD detector, whose output is fed into the distributional shift compensator to generate the rotational salience weight $\beta$. In the inference pipeline, given that the LoRA parameters $BA$ are proportional to the rotation angle $\theta_{R_{BA}}$ in the Cognitive Rotation Space $R_{BA}$. We control the rotation angle $\theta_{R_{BA}}$ amplitude by adjusting the scale of LoRA BABA, and use the weight $\beta$ to dynamically load the parameters that match the required unlearning degree.
@@ 3.1 Continuous Unlearning for LLM with LoRA
The continuous unlearning process of LLM inevitably leads to cumulative catastrophic utility loss (Gao et al. (2025)). The cumulative catastrophic utility loss manifests as a significant decline in both the LLM’s unlearning capability and its utility retention capability as the number of unlearning requests increases. At the parameter level, this is reflected in a shift of the parameters corresponding to previous unlearning requests when the model is trained on new ones. This requires the method to simultaneously achieve the continuous preservation of its historical unlearning knowledge and the original utility when handling the current unlearning request.
While existing approaches $o^{3}$ (Gao et al. (2025)) rely on orthogonal constraints to enforce perpendicularity between parameters of AA ($W=BA$). However, this constraint suffers from inherent limitations: the introduction of parameter $B$ undermines its effectiveness, and its capability further diminishes as the number of unlearning requests accumulates, making it inadequate for continuous unlearning scenarios. Moreover, the $o^{3}$ empirically assigns weights to LoRA parameters to represent the degree of unlearning, which is a heuristic design that lacks theoretical grounding and leads to poor interpretability.
We proposed the RCU to address these challenges. We develop a mathematically-derived approach with enhanced interpretability for quantifying unlearning process. By constructing a cognitive rotation space wherein LLM parameter updates are formulated as a novel rotational transformation, the RCU transforms continuous unlearning into rotations in a high-dimensional parameter space, thereby converting uncontrollable parameter shift into controlled angular rotations and ultimately alleviating cumulative catastrophic utility loss.
Our analysis in Figure 2 (a)(b) revealed the relationship between different weighting coefficients $\beta$ and unlearning updates, demonstrating that the extent of unlearning intensifies with increasing values of the $\beta$.
@@
$W^{\prime}\leftarrow W+BA.$ (3)
LoRA updates inevitably lead to uncontrollable parameter shift. To address this problem, we draw inspiration from Lie group (Gallier (2001)) to introduce a new update paradigm. This paradigm redefines parameter updates as rotations within the original parameter space. As rotations are the rigid transformation (Daniele (2001)), the update via the $BA$ matrices solely governs the change in the rotation angle. Consequently, the unlearning process is directly driven by changes in this angle, enabling us to use the rotation angle to precisely quantify the degree of unlearning in LLMs.
@@
However, every element $C$ in the Lie algebra $\mathfrak{so}\left(n\right)$ can be mapped to the $R$ in the Lie group $\mathit{SO}\left(n\right)$ by the exponential map $exp\left(C\right)$. From this we obtain the cognitive rotation space $R$:
@@
In addition, due to the influence of factors such as the learning rate, $BA<<I$ (as summarized in Table 4 and Figure 6 of Appendix A.6, the specifications for the $BA$ matrix parameters are detailed there.). Thereby, there exists a cognitive rotation space $R_{BA}\approx I+BA$.
@@
We hope that $\bigtriangleup R_{t}=I+B_{t}A_{t}-B_{t-1}A_{t-1}$ and $R_{t-1}=I+B_{t-1}A_{t-1}$ are perpendicular to each other, then $B_{t}A_{t}-B_{t-1}A_{t-1}$ and $B_{t-1}A_{t-1}$ will also be perpendicular to each other from Equation 5 and Theorem 2. The orthogonal rotation axes loss are as follows:
$\mathcal{L}_{o}=\left\|\left(W_{t}-W_{t-1}\right)\cdot W_{t-1}\right\|_{F}^{2}=\left\|\left(B_{t}A_{t}-B_{t-1}A_{t-1}\right)\cdot\left(B_{t-1}A_{t-1}\right)\right\|_{F}^{2},$ (11)
@@
here,we set $\lambda_{1}=0.1$$\lambda_{2}=0.1$ and $\lambda_{3}=1$ on the ScienceQA dataset. We set $\lambda_{1}=0.01$$\lambda_{2}=0.5$ and $\lambda_{3}=1$ on the TOFU dataset.
@@ 4.2 Experimental Setup
Table 1: Performance Comparison between our method and other baselines when continually unlearning TOFU-forget01, -forget05, and -forget10 in Fictitious Knowledge Generation. The * represents the results we achieved in our own experimental environment.
@@ 4.3 Experimental Results
Question Answering. The results of our method are shown in appendix A.8 and Figure 3. We compared base, PO, SOPO, $o^{3}$ and our results. Some detailed results can be found in the appendix. Our method achieved the same results as the base method (C.QA. and O.QA.). The results of the R.D. were also very similar to those of the base method. However, our method performed much better than the others in the S.U. and D.U. indicators. Compared with the most advanced method $o^{3}$, the S.U. decreased by 9.68%, 4.02%, 4.05%, 7.58% and 10.82% respectively on five requests. The D.U. compared to the most advanced method $o^{3}$, on average, decreased by 17.67% across five requests. In addition, the number of training parameters of our method is much smaller than that of $o^{3}$.
Fictitious Knowledge Generation. Table 1 presents the experimental results of our method on the TOFU dataset. Based on the results, we found that our method can effectively enhance the LLMs’ ability to unlearning (D.U. and S.U.). Compared with the currently best continuous unlearning method $o^{3}$, the result on the S.U. decreased by 6.29%, 9.99% and 6.06%. the result on the D.U. decreased by 2.17%, 7.67% and 6.06%. Furthermore, in terms of the stability of the LLMs’ utility, the performance level of our method is comparable to the current best results. We observed that as the number of unlearning requests increased, the unlearning accuracy on the test set (D.U.) of most methods showed significant fluctuations or increasing. This phenomenon indicates that catastrophic utility loss continues to accumulate, and the LLMs’ stability is affected. In contrast, our method not only effectively mitigates such utility losses but also significantly improves the performance stability of LLMs during the continuous unlearning process.
@@
Furthermore, we follow the Unlearning-Utility Ratio ($U^{2}R$) in $o^{3}$.
$U^{2}R=\frac{\mathrm{Acc}_{\mathrm{S}.\mathrm{U}.}^{0}+\mathrm{Acc}_{\mathrm{D}.\mathrm{U}.}^{0}-\mathrm{Acc}_{\mathrm{S}.\mathrm{U}.}^{T}-\mathrm{Acc}_{\mathrm{D}.\mathrm{U}.}^{T}}{\mathrm{Acc}_{\mathrm{R}.\mathrm{D}.}^{0}+\mathrm{Acc}_{\mathrm{U}.1.}^{0}+\mathrm{Acc}_{\mathrm{U}.2.}^{0}-\mathrm{Acc}_{\mathrm{R}.\mathrm{D}.}^{T}-\mathrm{Acc}_{\mathrm{U}.1.}^{T}-\mathrm{Acc}_{\mathrm{U}.2.}^{T}},$ (19)
where the Acc denotes Accuracy, the U.1 represents C.QA or R.A., the U.2 represents O.QA or W.F., and the T denotes the unlearning request. The result on the TOFU dataset is shown in Figure 2 (c). Our method has demonstrated significant advantages in both unlearning and utility retention. Furthermore, the results on the ScienceQA dataset can be found in the Appendix A.2.
Figure 2: Experimental results on the ScienceQA dataset. (a) The relationship between the $\beta$ ($\beta$ is proportional to the rotation angle) and unlearning processes (S.U.). (b) The relationship between the $\beta$ ($\beta$ is proportional to the rotation angle) and unlearning processes (D.U.). (c) The results of $U^{2}R$ on the TOFU dataset.
4.4 The Relationship between Rotation and Unlearning
In Figure 2 (a) and Figure 2 (b), we have presented the changes about $\beta$, along with the experimental results for each unlearning request. Based on the results on the ScienceQA dataset, we can observe that when $\beta$ is within the range of 0.15 to 0.45, the model rapidly undergoes the process of unlearning. However, when $\beta$ is less than 0.15, the model does not forget the knowledge. And when $\beta$ is greater than 0.45, the model has achieved complete forgetting and no longer continues to unlearning. The above experimental results indicate that the unlearning process of RCU is achieved within a very small rotation angle change range in the cognitive rotation space. Therefore, when designing the OOD detector and the distributional shift compensator, we must also ensure that the rotational salience weights they output have a relatively concentrated distribution to match the above characteristics. The $\beta$ changes regarding TOFU dataset are presented in the appendix.
@@ 5 Conclusion
Existing LLM unlearning methods are vulnerable not only to cumulative catastrophic utility loss from continuous unlearning requests, but also to a fundamental limitation in practicality owing to their heavy reliance on the retained dataset. To overcome this limitation, we introduce RCU, a retain-free approach that formulates LoRA updates as rotations within a specially constructed cognitive rotation space. This formulation makes the rotational angle updated by LoRA the sole variable, which is strongly correlated with the degree of unlearning. Consequently, the unlearning process can be directly quantified by corresponding changes in the rotational angle. Furthermore, by analyzing the update dynamics during unlearning, we propose a rotational salience weight to achieve precise and continuous control over the unlearning process. Our method is supported by theoretical guarantees, and we rigorously establish its efficacy through a mathematical analysis. Extensive experiments demonstrate that RCU achieves superior unlearning effectiveness while maintaining model utility on multiple datasets.
@@ Appendix A Appendix
• A.4: The experimental results on the TOFU dataset showing the relationship between $\beta$ and the unlearning process.
• A.5: The detailed calculation steps of the OOD Score.
• A.6: An in-depth study on why our method is effective.
• A.7: Experimental results regarding the hyperparameter $\lambda$
• A.8: Computation overhead analysis.
• A.9: The comparison results of the Truth Ratio on the TOFU dataset.
@@
$exp\left(kC\right)$ $=exp\left(C^{\prime}\right)$ (23)
$=Q\cdot e^{\left(diag\left(0,...,0,B\left(k\theta_{1}\right),...,B\left(k\theta_{m}\right)\right)\right)}\cdot Q^{T}$
$=Q\cdot diag\left(1,...,1.B\left(k\theta_{1}\right),...,B\left(k\theta_{m}\right)\right)\cdot Q^{T}.$
@@ A.2 More Experimental Results
The detailed results of our work on the ScienceQA dataset are shown in Table 3. Based on the results, we can observe that compared to $o^{3}$, our method is more stable and has a better ability to resist cumulative catastrophic utility loss. Here, the $U^{2}R$ of our method is 65.79, and the $U^{2}R$ of the other method is 48.46.
Table 3: Performance Comparison between our method and other baselines when continually unlearning biology, physics, chemistry, economics and earth-science in Fictitious Knowledge Generation. The unlearning effectiveness is measured by the generation accuracy of the unlearning train data and unlearning test data denoted as S.U. and D.U., CommonsenseQA (C.QA.), OpenbookQA (O.QA.) respectively. Utility preservation is evaluated by the generation accuracy of Retained Distribution (R.D.). The * represents the results we achieved in our own experimental environment.
@@
The specific results are shown in Figure 4.
Figure 4: Experimental results on the ScienceQA dataset. (a) The results of C.QA.. (b)The results of O.QA..
@@ A.3 Large Language Model Usage Declaration
A.4 Hyperparameter analysis on the TOFU dataset
@@ A.5 Calculation of OOD Score
A.6 Research on Our Method
Unlearning Request Param $B$ Param AA
biology $10^{-6}$ $10^{-5}$
physics $10^{-5}$ $10^{-5}$
chemistry $10^{-5}$ $10^{-5}$
economics $10^{-6}$ $10^{-5}$
earth-science $10^{-6}$ $10^{-5}$
Furthermore, Figure 6 (a)(b) present the evolution of parameters $\Theta_{A}$ and $\Theta_{B}$ throughout the experiment. The results indicate that the magnitudes of both $\Theta_{A}$ and $\Theta_{B}$ remain significantly smaller than those of parameter I during the entire training process.
Figure 6: The results of 5 continuous unlearning processes on the ScienceQA dataset. (a) The changes in the $\Theta_{A}$. (b) The changes in the $\Theta_{B}$.
A.7 Hyperparameter Analysis
We conducted a comprehensive hyperparameter analysis on $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ across two datasets. The results, presented in Figure 7 and Figure 8, demonstrate that our chosen hyperparameter configuration consistently achieves optimal performance.
Figure 7: Experimental results on the ScienceQA dataset. The left column shows the results of S.U., the middle column shows the results of D.U., and the right column shows the results of R.D.. (a) (b) (c)The result of the hyperparameter $\lambda_{1}$. (d) (e) (f)The result of the hyperparameter $\lambda_{2}$(S.U.). (g) (h) (i)The result of the hyperparameter $\lambda_{3}$(S.U.).
A.8 Computation overhead analysis.
During the reasoning process, the computational cost for LLM to update using the lora update paradigm proposed by us is 45.10 GFLOPs, the computational cost for OOD Detection is 709 MFLOPs, and the computational cost for the Distributional Shift Compensator is 13255 MFLOPs. The total computational cost is 59.064 GFLOPs. In addition, our method requires less storage space. The storage space required by lora in $o^{3}$ is 39MB, while our method only requires 16MB, reducing the additional storage space requirement by 58%.
A.9 The comparison results of the Evaluation Indicators on the TOFU dataset.
We choose the Truth Ratio and ROUGE-L as the indicators for evaluating the utility on the TOFU dataset.
The results of the Truth Ratio (Maini et al. (2024)) are presented in Table 5. From these results, it can be seen that our method is superior to the comparison methods. The formula for the Truth Ratio is as shown in Equation 30.
$R_{\text{truth }}=\frac{\frac{1}{\mathcal{A}_{\text{pert }}}\sum_{\hat{a}\in\mathcal{A}_{\text{pert }}}P(\hat{a}\mid q)^{1/|\hat{a}|}}{P(\tilde{a}\mid q)^{1/|\tilde{a}|}}$, (30)
[…]
Table 5: The results here are all the Truth Ratio (Maini et al. (2024)) corresponding to the aforementioned indicators.
[…]
$o^{3}$ 0.74 0.66 0.57 0.54 1.54 0.65 0.65 0.57 0.54 1.54 0.66 0.65 0.57 0.54 1.54
Ours 1.00 1.11 0.56 0.54 1.22 0.99 0.97 0.56 0.54 1.22 0.97 1.01 0.56 0.54 1.22
We compute the ROUGE-L recall score (Maini et al. (2024)), which acts as a surrogate for accuracy on the question answering task, as it accounts for the output phrasing to be slightly different than the ground truth. The results of the ROUGE-L (Maini et al. (2024)) are presented in Table 6. From these results, it can be seen that our method is superior to the comparison methods.
Table 6: The results here are all the ROUGE-L (Maini et al. (2024)) corresponding to the aforementioned indicators.
[…]
$o^{3}$ 0.0771 0.0627 0.9675 0.9330 0.8960 0.1939 0.1255 0.9677 0.9330 0.8960 0.1843 0.5189 0.9675 0.9330 0.8960
Ours 0.0425 0.0365 0.9675 0.9330 0.9083 0.1333 0.1044 0.9683 0.9330 0.9083 0.1322 0.1062 0.9683 0.9330 0.9083
Figure 8: Experimental results on the TOFU dataset. The left column shows the results of S.U., the middle column shows the results of D.U., and the right column shows the results of R.D.. (a) (b) (c)The result of the hyperparameter $\lambda_{1}$. (d) (e) (f)The result of the hyperparameter $\lambda_{2}$(S.U.). (g) (h) (i)The result of the hyperparameter $\lambda_{3}$(S.U.).